The area of a garden can be found using the equation
A=2n^2-21n, where n>10
What is the value of n when the area is 176 square units

Answer by Cromlix(4381)

Hi there,
2n^2 - 21n = 176
2n^2 - 21n - 176 = 0
(2n + 11)(n - 16) = 0
2n + 11 = 0
n = -5.5 (Disregard as -ve)
n - 16 = 0
n = 16
Hope this helps :-)